OpenStudy (anonymous):

for the function f(x)=x^3+2x, approximate the area between the curve and the x-axis over the interval [2, 0] using four equal subintervals and a) a left-hand Riemann sum b) a right-hand Riemann sum

OpenStudy (anonymous):

ok we can do this but first of all there is no such interval is [2,0]

OpenStudy (anonymous):

i assume it is [0,2]

OpenStudy (anonymous):

lol, typo :]

OpenStudy (anonymous):

ok so we divide the interval into 4 equal parts. lets do a) first

OpenStudy (anonymous):

[0,2] has length 2, so dividing it into 4 equal parts means they will have length 1/2

OpenStudy (anonymous):

clear yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait, length or width?

OpenStudy (anonymous):

so we get 0, .5, 1, 1.5, 2 as the numbers for the partition

OpenStudy (anonymous):

width i suppose since we are on the x - axis so don't let me confuse you

OpenStudy (anonymous):

so now it is easy. you calculate base times height for each of the 4 rectangles. in part a) we use the left hand endpoints, which are 0 , .5, 1, 1.5 and compute \[f(0)\times \frac{1}{2}+f(.5)\times \frac{1}{2}+f(1)\times \frac{1}{2} +f(1.5)\times \frac{1}{2}\] \[=(f(0)+f(.5)+f(1)+f(1.5))\frac{1}{2}\]

OpenStudy (anonymous):

on part b) use the right hand endpoints to get \[(f(.5)+f(1)+f(1.5)+f(2))\times \frac{1}{2}\]

OpenStudy (anonymous):

i would use a calculator myself. or maple if i could access it from this operating system. you want me to check the first answer?

OpenStudy (anonymous):

5.25 for the first one?

OpenStudy (anonymous):

and 11.25 for the second?

OpenStudy (anonymous):

yeah i got 5.25 for the first one too